# Algorithmic design of “smart canopy” maize architectures that maximize light use efficiency

### v1

@@ Abstract
keywords:
Maize “smart canopy” design , Photosynthetically active radiation (PAR) , evolutionary optimization
[…]
[inst1]organization=Department of Mechanical Engineering,addressline=Iowa State University, city=Ames, postcode=50011, state=Iowa, country=USA
[…]
@@ 1 Introduction
Maize (Zea mays L.) is a cornerstone of global agriculture, with annual production exceeding 1.1 billion metric tons and serving as a primary source of food, feed, fuel, and industrial products erenstein2022global, nyirenda2021delving. As the world’s population continues to grow and climate pressures intensify, maximizing maize yields on existing cropland is paramount. Increasing agricultural production on the same land base also reduces pressure to convert natural ecosystems, such as the Amazon rainforest, into new agricultural areas, thereby helping preserve biodiversity and carbon-rich landscapes. Over the past century, one of the most successful strategies for yield improvement has been increasing planting density: in the United States, average planting densities rose from roughly 30,000 plants ha-1 in the mid-20th century to over 80,000 plants ha-1 today assefa2018analysis, yan2024photosynthetic, kalogeropoulos2024historical, duvick2005contribution. This intensification was enabled by selecting at high plant densities for higher-yielding hybrids. Selection for higher yield under these conditions was associated with indirect selection of hybrids with more upright leaves and compact canopies that mitigate self-shading and improve light capture in dense stands. Yet, further gains via increased density alone are increasingly constrained by inter-plant competition and mutual shading, which deprive lower canopy leaves of light and cap whole-canopy photosynthesis mansfield2014survey, saleem2025accessing.
@@
The concept of an idealized crop architecture, or ideotype, was first articulated by Donald in 1968 as a set of trait values optimized for a given environment and management system donald1968breeding, mock1975ideotype, carbajal2024role, gauffreteau2018using, sedgley1991appraisal. In grain crops, the so-called “smart canopy” ideotype balances light distribution such that upper leaves are erect, allowing light to penetrate, while lower leaves are more horizontal to intercept the attenuated irradiance deep in the canopy ort2015redesigning. In maize, studies have borne out this paradigm: genotypes exhibiting a vertical gradient of leaf angles demonstrate enhanced light penetration and yield under high densities jiang2025leaf, pepper1977leaf, han2024molecular. Breeding for higher yield over the past several decades has also increased the prevalence of erect upper leaves, a trait consistently associated with improved light distribution and density tolerance and noted as a contributing factor to historical yield gains duvick2005contribution. However, recent analyses indicate that the rate of genetic gain in vertical leaf angle itself is slowing elli2023maize, suggesting that further improvement of this trait may be reaching its practical limits under current agronomic planting densities. This plateau underscores the need to identify and optimize additional canopy traits beyond upright leaves at the top to further improve light distribution and sustain yield gains in dense plantings.
@@
Traditional breeding, however, relies on incremental selection constrained by the ’breeder’s eye’ and the logistical bottlenecks of field trials hossain2022maize, murchie2022casting. Consequently, the sheer dimensionality of canopy design—encompassing leaf length, width, vertical, and azimuthal angles—creates a combinatorial landscape too vast for empirical exploration murchie2022casting, leaving potentially beneficial architectural solutions undiscovered. Moreover, many functional-structural plant models vos2007functional, vos2010functional impose fixed allometries or narrow trait ranges, limiting their ability to propose truly novel architectures liu2025maize2035, carbajal2024role, he2021modeling, gage2019field. Here, we present a computational ideotype-design framework that systematically explores this vast trait space and uncovers high-performing canopy architectures beyond the sampling of existing germplasm. Through this process, we identify an emergent ideotype—a coordinated combination of architectural traits—that maximizes canopy-scale light interception under dense planting conditions. Our approach comprises three integrated components:
@@
  • Physics-based light interception simulation: Using a radiative-transfer model bailey2018reverse, bailey2019helios, we compute daily photosynthetically active radiation (PAR) mccree1981photosynthetically intercepted by each virtual canopy under realistic solar trajectories and field-density scenarios (fig 1E). Our simulation workflow follows a previously documented, validated pipeline; optical parameters and multiple-scattering settings are detailed in the SI and prior work saleem2025accessing.
@@
  • Evolutionary algorithm (specifically, genetic algorithm (GA)): Evolutionary algorithms systematically improve a set of virtual canopies by recombining and mutating trait vectors, then selecting the highest-performing individuals over many generations. This search heuristic excels in navigating complex, multimodal design landscapes where gradient-based methods falter holland1992genetic, whitley1994genetic. Overview of the framework is shown in Fig 1 A-H.
@@
Figure 1: Computational framework for canopy optimization. (A) A parametric 3D maize plant model was constructed with leaf length, width, leaf vertical angle, and azimuth as controllable traits. (B) Virtual fields were generated by replicating these models at fixed row and plant spacing. (C) Reverse ray-tracing simulations quantified diurnal photosynthetically active radiation (PAR) absorption with multiple scattering. (D) Daily intercepted PAR profiles highlight improved interception dynamics compared with the baseline. (E) An evolutionary algorithm (detailed in Fig. 2) iteratively modified plant traits, evaluated PAR interception, and selected beneficial architectures. (F) Convergence plots demonstrate progressive improvement in canopy-level PAR interception across generations. (G) The resulting optimized canopy architecture exhibited enhanced light capture. (H) Virtual field of the optimized ideotype.
@@
By treating key architectural features as free variables, the opimization revealed emergent strategies that a priori were not obvious. In addition to the expected vertical stratification of leaf angles consistent with the smart canopy, our optimized ideotype exhibited a distinctive radially tiled arrangement: successive leaves were azimuthally offset, reducing self- and neighbor-shading within and between plants. As detailed in the Results and SI (Sections S4–S5), the qualitative canopy organization (vertical stratification of angles and widths, tiled azimuths) was preserved across multiple maize-growing latitudes and six spacing regimes, indicating robustness to solar geometry and density.
@@
Beyond uncovering novel trait combinations, our framework offers considerable practical advantages: (i) Scalability: We can rapidly evaluate thousands of ideotypes across a range of planting densities, row spacings, and geographic latitudes, enabling region-specific customization; (ii) Extensibility: The open-source pipeline can incorporate additional constraints, such as currently accessible allometric constraints, or additional traits like mechanical stability or water-use efficiency, or be adapted to other crop species, or other planting arrangements; and (iii) Translatability: Advances in gene-editing technologies like CRISPR/Cas9 now make it feasible to target specific architectural genes independently, surmounting the challenges of genetic linkage and pleiotropy that hinder conventional breeding wang2022analysis, wang2022crispr, liu2020high. In practice, however, implementing these modifications still requires identifying the genes that control each trait and determining how specific edits will produce the desired phenotypic effect. By pinpointing precise trait values, our computational ideotype design can directly inform molecular breeding and genome-editing strategies.
@@
Figure 2: Evolutionary Algorithm framework (specifically, a Genetic Algorithm) for optimizing maize architecture. The process begins with an initial population of maize plant architectures, each defined by key parameters (leaf length, width, angle, and azimuth orientation). The EA iteratively evaluates plant designs based on their light capture efficiency, applies selection, crossover, and mutation operations to generate new candidate solutions, and refines the population over successive generations until an optimal maize architecture emerges.
@@
We developed a parametric 3D maize canopy model based on Non-Uniform Rational B-Spline (NURBS) surfaces, incorporating realistic leaf-level structural parameters (hereafter referred to as traits) piegl2012nurbs, hadadi2025procedural. Each virtual plant consisted of a fixed number of leaves (ten per plant, consistent with typical maize architecture) arranged along a central stalk. Each leaf was represented as a NURBS surface, defined by control points that describe its curvature and orientation in 3D space. Four independent traits – leaf length, maximum width, inclination angle (from the stem), and azimuth (rotation around the vertical axis) – were parameterized for each leaf. Trait bounds were derived from empirical field measurements across diverse maize genotypes in high-density plots (see Section S1 in the SI Appendix), and from leaf definitions used in a curated, leaf-annotated 3D maize resource, ensuring that all simulated architectures remained biologically feasible and mesh-interoperable MaizeField3D. These traits capture the dominant structural determinants of canopy light interception while keeping the model computationally efficient for large-scale virtual experiments.
To construct the virtual canopy, a single 3D plant mesh was exported from the NURBS model and replicated in a $6\times 6$ grid. Plants were positioned according to a row spacing of 30 inches (0.76 m) and a within-row plant spacing of 6 inches (0.15 m), matching typical high-density planting configurations. To mimic realistic per-plant variability, we did not tile identical copies of a candidate: for each plant in the virtual field we independently applied a stochastic azimuthal jitter of $\pm 10^{\circ}$ about the vertical axis to capture orientation variability. The individual meshes were combined into a composite field geometry. Periodic boundary conditions were applied to minimize edge effects, effectively simulating an infinite canopy. The resulting field-level mesh was exported as an OBJ file for light interception modeling in Helios bailey2019helios.
Simulations targeted the R1 (silking) stage, approximately 60 days after planting abendroth2011corn. At this stage, the maize canopy architecture is effectively fixed: all leaves are fully expanded, internode elongation has ceased, and leaf angles and orientations remain stable abendroth2011corn. Therefore, we assume a static canopy for modeling light interception. Additionally, from the R1 stage onward, a substantial portion of photosynthates is partitioned toward the kernels, making intercepted PAR a meaningful proxy for yield potential tollenaar1999physiology, andrade2002yield, lizaso2017modeling.
@@
Intercepted PAR was simulated using the radiation module of Helios bailey2019helios (v2.1.0). The model computes dynamic solar position and mutual shading at mesh-level geometric precision. Simulations (for the results in the main text) assumed a representative clear-sky day in Ames, Iowa (42.0N, 93.6W) on August 7, spanning 07:00–20:00 local time at hourly resolution. Both direct and diffuse components of solar irradiance were modeled, assuming maize leaf reflectance =0.05=0.05 and transmittance =0.05=0.05 following Earl and Tollenaar earl1997maize. Instantaneous absorbed PAR fluxes (W m-2) were integrated over the daylight period and converted to daily energy per ground area, expressed as MWh acre-1 day-1. This daily canopy-integrated absorbed PAR served as the fitness metric for the evolutionary algorithm (EA) optimization described below.
All Helios configuration parameters, including optical properties and atmospheric inputs, are provided in Section S2 of the SI Appendix. The configuration follows a previously documented, field-referenced pipeline for estimating canopy PAR saleem2025accessing.
@@
To identify canopy architectures that maximize light capture, we employed an evolutionary algorithm (EA) lambora2019genetic. Each candidate plant was represented as a “chromosome” comprising 40 “gene”, corresponding to four parameters (leaf length, width, inclination, and azimuth) for each of ten leaves. The EA initialized a population of 100 diverse canopy architectures and evolved them over 200 generations to maximize the total daily absorbed photosynthetically active radiation (PAR) per acre. The optimization workflow is summarized in Fig. 2.
@@
Results
Building on the computational framework described above, we evaluated how the evolutionary algorithm reshaped canopy architecture and light interception performance under high‐density planting. Simulations were conducted using a parameterized 3D maize model comprising ten leaves per plant, each independently defined by length, width, insertion angle, and azimuthal orientation (Fig. 1A). Virtual plants were arranged in a field layout representative of modern high-density conditions, with 30-inch (0.76 m) row spacing and 6-inch (0.15 m) plant spacing, corresponding to approximately $87{,}700~\text{plants}~\text{ha}^{-1}$, and periodic boundary conditions were applied to minimize edge effects. To mimic realistic per‐plant architectual variability, we do not tile identical copies of a candidate: instead, for each plant in the virtual field we independently applied a stochastic azimuthal jitter of $\pm 10^{\circ}$ about the vertical axis to capture orientation variability. Trait ranges were constrained by field‐measured values from diverse maize genotypes to ensure biological realism MaizeField3D (also see SI for data on ranges). All simulations represented the R1 (silking) stage, when canopy architecture is fixed and photosynthetically active radiation (PAR) interception closely approximates yield potential tollenaar1999physiology, lizaso2017modeling.
@@ Results
mThe evolutionary optimization algorithm successfully evolved canopy architectures that substantially enhanced light interception relative to the baseline virtual canopy derived from field‐measured hybrids saleem2025accessing. By iteratively adjusting key morphological traits—including vertical leaf angle, width, length, and azimuthal leaf angle—the optimization explored a wide range of canopy architectures and converged on an ideotype that maximized canopy‐scale PAR interception under dense planting conditions (Fig. 1A).
Across three independent optimization runs with different random initializations, the algorithm consistently converged to a canopy configuration intercepting 56.49 MWh PAR acre-1 day-1, representing a 16% improvement over field hybrid with one of the highest PAR interceptions - 48.72 MWh PAR acre-1 day-1. Fitness trajectories stabilized after approximately 80 generations, indicating convergence of the evolutionary search (Fig. 1F). The similarity among final ideotypes across runs suggests robustness and repeatability of the optimization process.
The following sub-sections characterize the structural and functional features of these optimized ideotypes, highlighting how architectural modifications improved light distribution throughout the canopy and their broader implications for breeding and crop design.
@@
Daily PAR interception profiles revealed that optimized canopy architectures consistently outperformed the baseline across all hours of the day (Fig. 1D). Improvements were greatest during mid‐morning and late afternoon when conventional canopies suffer from self‐shading; the optimized architecture maintained higher PAR penetration throughout the canopy. Integrating PAR intercepted over time yielded the 16% daily improvement noted above. Simulations across a range of latitudes (SI, Section S4) showed similar gains, indicating that the ideotype is not over‐fit to a single environment. We further evaluated the optimization across a range of planting densities and row spacings; detailed results are provided in SI, Section S5. Radiative‐transfer configuration followed a documented Helios pipeline (reverse ray tracing with multiple scattering), previously shown to agree with field PAR interception measurements, providing additional confidence in the simulated gains bailey2019helios, saleem2025accessing.
@@
The optimized ideotype featured clear vertical stratification of leaf traits. Leaf angles were measured relative to the stalk, such that smaller angles indicate more erect leaves (0 = parallel to the stalk, 90 = horizontal). Upper canopy leaves exhibited smaller inclination angles (20–50), effectively minimizing top-layer shading, whereas lower canopy leaves displayed higher inclination angles (60–80), improving light interception in shaded regions. A similar vertical gradient was observed in leaf width, where upper leaves were narrower and lower leaves became progressively wider, enabling efficient capture of the residual sunlight that penetrates deeper into the canopy (Fig. 3). This configuration promotes efficient light redistribution by allowing incoming radiation to penetrate deeper into the canopy and be absorbed more uniformly across layers. Such trait stratification emerged consistently across independently evolved solutions, suggesting a convergent optimization strategy driven by canopy-scale light dynamics.
@@
We also examined how leaves were azimuthally arranged around the stalk, a trait that influences both self-shading and interplant light competition. Maize naturally exhibits an alternate (distichous) phyllotaxy, with leaves emerging in two opposing ranks along the stem. In field-grown maize, this inherent pattern typically manifests as one of three canopy-level azimuthal orientations: off-row parallel (perpendicular to the planting row), which minimizes mutual shading; on-row parallel, where leaves align with the row direction; or random, where entire individual plants vary in their overall orientation relative to the planting rows. saleem2025accessing, zhou2024genetic
In contrast to the strict alternate phyllotaxy typical of maize, our EA optimization converged upon a tiled or quasi-radial arrangement of leaf azimuths around the stalk (Fig. 4). This deviation suggests that while the canonical distichous structure may optimize light capture for isolated plants, it generates excessive mutual shading in dense monocultures. The emergent radial tiling represents a ’community-optimal’ strategy that minimizes self- and neighbor competition.
@@
The optimized ideotype exhibited architectural features consistent with both experimentally validated and commercially successful maize hybrids, underscoring the practical relevance of computational design. Our results on leaf angles closely resemble the smart-canopy architecture characterized by Tian et al. tian2024maize in lac1-based F1 hybrids, which display upright upper, moderately inclined middle, and flatter lower leaves—traits that improved light penetration and boosted grain yield by 10–25% under high planting densities. The resemblance between our computationally derived canopy and this experimentally validated smart-canopy structure reinforces the physiological credibility and predictive capacity of the optimization framework. A similar pattern occurs in elite commercial hybrids such as P1151AM™ and P1311AMXT™, where erect upper leaves enhance light transmission and density tolerance [PioneerLeafAngle2023].
@@
To verify that these architectural patterns were mechanistically responsible for improved light capture, we performed trait-restricted optimization experiments in which the EA was restricted to optimize only a subset of parameters while all others were clamped to baseline values. When only leaf angles were free, the algorithm consistently produced ideotypes with erect upper leaves, achieving a modest gain in PAR interception (40.72 $\rightarrow$ 43.00 MWh acre-1 day-1). Optimization of azimuths alone yielded a tiled leaf orientation around the stalk, with a comparable improvement (40.72 $\rightarrow$ 42.99 MWh acre-1 day-1). Allowing angles and widths together led to erect, narrower upper leaves, substantially increasing canopy interception (40.72 $\rightarrow$ 49.47 MWh acre-1 day-1). Finally, when all traits were optimized jointly, the algorithm integrated these features synergistically to reach the highest performance (56.49 MWh acre-1 day-1).
These trait-restricted analyses demonstrate that the key architectural signatures identified in the full optimization (erect angles, tiled azimuths, and thinner upper leaves) emerge independently when traits are optimized in isolation, and that their combination yields additive and synergistic benefits to canopy light capture (Table 1).

### v2

# Towards smart canopies: Algorithmic design of maize canopy architectures that maximize light use efficiency

@@ Abstract
Keywords: Maize “smart canopy” design; Photosynthetically active radiation (PAR); evolutionary optimization.
@@ 1 Introduction
Maize (Zea mays L.) is a cornerstone of global agriculture, with annual production exceeding 1.1 billion metric tons and serving as a primary source of food, feed, fuel, and industrial products [erenstein2022global, nyirenda2021delving]. As the world’s population continues to grow and climate pressures intensify, maximizing maize yields on existing cropland is paramount. Increasing agricultural production on the same land base also reduces pressure to convert natural ecosystems, such as the Amazon rainforest, into agricultural production, thereby helping to preserve biodiversity and carbon-rich landscapes. Over the past century, one of the most successful strategies for yield improvement has been increasing planting density: in the United States, average planting densities rose from roughly 30,000 plants ha-1 in the mid-20th century to over 80,000 plants ha-1 today [assefa2018analysis, yan2024photosynthetic, kalogeropoulos2024historical, duvick2005contribution]. This intensification was enabled by selecting at high plant densities for higher-yielding hybrids. Selection for higher yield under these conditions was associated with indirect selection of hybrids with more upright leaves and compact canopies that mitigate self-shading and improve light capture in dense stands. Yet, further gains via increased density alone are increasingly constrained by inter-plant competition and mutual shading, which deprive lower canopy leaves of light and cap whole-canopy photosynthesis [mansfield2014survey, saleem2025accessing].
@@
The concept of an idealized crop architecture, or ideotype, was first articulated by Donald in 1968 as a set of trait values optimized for a given environment and management system [donald1968breeding, mock1975ideotype, carbajal2024role, gauffreteau2018using, sedgley1991appraisal]. In grain crops, the so-called “smart canopy” ideotype balances light distribution such that top leaves are erect, allowing light to penetrate, while lower leaves are more horizontal to intercept the attenuated irradiance deep in the canopy [ort2015redesigning]. In maize, studies have borne out this paradigm: genotypes exhibiting a vertical gradient of leaf angles demonstrate enhanced light penetration and yield under high densities [jiang2025leaf, pepper1977leaf, han2024molecular]. Breeding for higher yield over the past several decades has also increased the prevalence of erect upper leaves, a trait consistently associated with improved light distribution and density tolerance and noted as a contributing factor to historical yield gains [duvick2005contribution]. However, recent analyses indicate that the rate of genetic gain in vertical leaf angle itself is slowing [elli2023maize], suggesting that further improvement of this trait may be reaching its practical limits under current agronomic planting densities. This plateau underscores the need to identify and optimize additional canopy traits beyond upright leaves at the top to further improve light distribution and sustain yield gains in dense plantings.
@@
Traditional breeding, however, relies on incremental selection constrained by the ’breeder’s eye’ and the logistical bottlenecks of field trials [hossain2022maize, murchie2022casting]. Consequently, the sheer dimensionality of canopy design—encompassing leaf length, width, vertical, and azimuthal angles—creates a combinatorial landscape too vast for empirical exploration [murchie2022casting], leaving potentially beneficial architectural solutions undiscovered. Moreover, many functional-structural plant models [vos2007functional, vos2010functional] impose fixed allometries or narrow trait ranges, limiting their ability to propose truly novel architectures [liu2025maize2035, carbajal2024role, he2021modeling, gage2019field]. Here, we present a computational ideotype-design framework that systematically explores this vast trait space and uncovers high-performing canopy architectures beyond the sampling of existing germplasm. Through this process, we identify an emergent ideotype—a coordinated combination of architectural traits—that maximizes canopy-scale light interception under dense planting conditions. Our approach comprises three integrated components:
@@
  • Physics-based light interception simulation: Using a radiative-transfer model [bailey2018reverse, bailey2019helios], we compute daily photosynthetically active radiation (PAR) [mccree1981photosynthetically] intercepted by each virtual canopy under realistic solar trajectories and field-density scenarios (fig 1E). Our simulation workflow follows a previously documented, validated pipeline; optical parameters and multiple-scattering settings are detailed in the SI and prior work [saleem2025accessing].
@@
  • Evolutionary algorithm (specifically, genetic algorithm) for optimization: Evolutionary algorithms systematically improve a set of virtual canopies by recombining and mutating trait vectors, then selecting the highest-performing individuals over many generations. This search heuristic excels in navigating complex, multimodal design landscapes where gradient-based methods falter [holland1992genetic, whitley1994genetic]. Overview of the framework is shown in Fig 1 A-H.
@@
Figure 1: Computational framework for canopy optimization. (A) A parametric 3D maize plant model was constructed with leaf length, width, leaf vertical angle, and azimuthal angle as controllable traits. (B) Virtual fields were generated by replicating these models at fixed row and plant spacing. (C) Reverse ray-tracing simulations quantified diurnal photosynthetically active radiation (PAR) absorption with multiple scattering. (D) Daily intercepted PAR profiles highlight improved interception dynamics compared with the baseline. (E) An evolutionary algorithm (detailed in Fig. 2) iteratively modified plant traits, evaluated PAR interception, and selected beneficial architectures. (F) Convergence plots demonstrate progressive improvement in canopy-level PAR interception across generations. (G) The resulting optimized canopy architecture exhibited enhanced light capture. (H) Virtual field of the optimized ideotype.
@@
By treating key architectural features as free variables, the opimization revealed emergent strategies that a priori were not obvious. In addition to the expected vertical stratification of leaf angles consistent with the smart canopy, our optimized ideotype exhibited a distinctive radially tiled arrangement of leaves around the stalk: successive leaves were azimuthally offset, reducing self and neighbor-shading within and between plants. As detailed in the Results and SI (Sections S4–S5), the qualitative canopy organization (vertical stratification of angles and widths, tiled azimuths) was preserved across multiple maize-growing latitudes and six spacing regimes, indicating robustness to solar geometry and density.
@@
Beyond uncovering novel trait combinations, our framework offers considerable practical advantages: (i) Scalability: We can rapidly evaluate thousands of ideotypes across a range of planting densities, row spacings, and geographic latitudes, enabling region-specific customization; (ii) Extensibility: The open-source pipeline can incorporate additional constraints, such as currently accessible allometric constraints, or additional traits like mechanical stability or water-use efficiency, or be adapted to other crop species, or other planting arrangements; and (iii) Translatability: Advances in gene-editing technologies like CRISPR/Cas9 now make it feasible to target specific architectural genes independently, surmounting the challenges of genetic linkage and pleiotropy that hinder conventional breeding [wang2022analysis, wang2022crispr, liu2020high]. In practice, however, implementing these modifications still requires identifying the genes that control each trait and determining how specific edits will produce the desired phenotypic effect. By pinpointing precise trait values, our computational ideotype design can directly inform molecular breeding and genome-editing strategies.
@@
Figure 2: Evolutionary Algorithm framework (specifically, a Genetic Algorithm) for optimizing maize architecture. The process begins with an initial population of maize plant architectures, each defined by key parameters (leaf length, width, vertical angle, and azimuthal angle). The EA iteratively evaluates plant designs based on their light capture efficiency, applies selection, crossover, and mutation operations to generate new candidate solutions, and refines the population over successive generations until an optimal maize architecture emerges.
@@
We developed a parametric 3D maize canopy model based on Non-Uniform Rational B-Spline (NURBS) surfaces, incorporating realistic leaf-level structural parameters (hereafter referred to as traits) [piegl2012nurbs, hadadi2025procedural]. Each virtual plant consisted of a fixed number of leaves (ten per plant, consistent with typical maize architecture) arranged along a central stalk. Each leaf was represented as a NURBS surface, defined by control points that describe its curvature and orientation in 3D space. Four independent traits – leaf length, maximum width, inclination angle (from the stem), and azimuth (rotation around the vertical axis) – were parameterized for each leaf. Trait bounds were derived from empirical field measurements across diverse maize genotypes in high-density plots (see Section S1 in the SI Appendix), and from leaf definitions used in a curated, leaf-annotated 3D maize resource, ensuring that all simulated architectures remained biologically feasible and mesh-interoperable [MaizeField3D]. These traits capture the dominant structural determinants of canopy light interception while keeping the model computationally efficient for large-scale virtual experiments.
To construct the virtual canopy, a single 3D plant mesh was exported from the NURBS model and replicated in a $6\times 6$ grid. Plants were positioned according to a row spacing of 30 inches (0.76 m) and a within-row plant spacing of 6 inches (0.15 m), matching typical high-density planting configurations. To mimic realistic per-plant variability, we did not tile identical copies of a candidate: for each plant in the virtual field we independently applied a stochastic azimuthal jitter of $\pm 10^{\circ}$ about the vertical axis to capture orientation variability. The individual meshes were combined into a composite field geometry. Periodic boundary conditions were applied to minimize edge effects, effectively simulating an infinite canopy. The resulting field-level mesh was exported as an OBJ file for light interception modeling in Helios [bailey2019helios].
Simulations targeted the R1 (silking) stage, approximately 60 days after planting [abendroth2011corn]. At this stage, the maize canopy architecture is effectively fixed: all leaves are fully expanded, internode elongation has ceased, and leaf angles and orientations remain stable [abendroth2011corn]. Therefore, we assume a static canopy for modeling light interception. Additionally, from the R1 stage onward, a substantial portion of photosynthates is partitioned toward the kernels, making intercepted PAR a meaningful proxy for yield potential [tollenaar1999physiology, andrade2002yield, lizaso2017modeling].
@@
Intercepted PAR was simulated using the radiation module of Helios [bailey2019helios] (v2.1.0). The model computes dynamic solar position and mutual shading at mesh-level geometric precision. Simulations (for the results in the main text) assumed a representative clear-sky day in Ames, Iowa (42.0N, 93.6W) on August 7, spanning 07:00–20:00 local time at hourly resolution. Both direct and diffuse components of solar irradiance were modeled, assuming maize leaf reflectance =0.05=0.05 and transmittance =0.05=0.05 following Earl and Tollenaar [earl1997maize]. Instantaneous absorbed PAR fluxes (W m-2) were integrated over the daylight period and converted to daily energy per ground area, expressed as MWh acre-1 day-1. This daily canopy-integrated absorbed PAR served as the fitness metric for the evolutionary algorithm (EA) optimization described below.
All Helios configuration parameters, including optical properties and atmospheric inputs, are provided in Section S2 of the SI Appendix. The configuration follows a previously documented, field-referenced pipeline for estimating canopy PAR [saleem2025accessing].
@@
To identify canopy architectures that maximize light capture, we employed an evolutionary algorithm (EA) [lambora2019genetic]. Each candidate plant was represented as a “chromosome” comprising 40 “gene”, corresponding to four parameters (leaf length, width, inclination, and azimuth) for each of ten leaves. The EA initialized a population of 100 diverse canopy architectures and evolved them over 200 generations to maximize the total daily absorbed photosynthetically active radiation (PAR) per acre. The optimization workflow is summarized in Fig. 2.
@@
Building on the computational framework described above, we evaluated how optimization through evolutionary algorithm reshaped canopy architecture and light interception performance under high‐density planting. Simulations were conducted using a parameterized 3D maize model comprising ten leaves per plant, each independently defined by length, width, leaf vertical angle, and azimuthal angle (Fig. 1A). Virtual plants were arranged in a field layout representative of modern high-density conditions, with 30-inch (0.76 m) row spacing and 6-inch (0.15 m) plant spacing, corresponding to approximately $87{,}700~\text{plants}~\text{ha}^{-1}$, and periodic boundary conditions were applied to minimize edge effects. To mimic realistic per‐plant architectual variability, we do not tile identical copies of a candidate: instead, for each plant in the virtual field we independently applied a stochastic azimuthal rotation of $\pm 10^{\circ}$ about the vertical axis to capture orientation variability. Trait ranges were constrained by field‐measured values from diverse maize genotypes to ensure biological realism [MaizeField3D] (also see SI for data on ranges). All simulations represented the R1 (silking) stage, when canopy architecture is fixed and photosynthetically active radiation (PAR) interception closely approximates yield potential [tollenaar1999physiology, lizaso2017modeling].
@@
The evolutionary optimization algorithm successfully evolved canopy architectures that substantially enhanced light interception relative to the baseline virtual canopy derived from field‐measured genotypes [saleem2025accessing]. By iteratively adjusting key morphological traits—including vertical leaf angle, width, length, and azimuthal leaf angle—the optimization explored a wide range of canopy architectures and converged on an ideotype that maximized canopy‐scale PAR interception under dense planting conditions (Fig. 1A).
Across three independent optimization runs with different random initializations, the algorithm consistently converged to a canopy configuration intercepting 56.49 MWh PAR acre-1 day-1, representing  13% improvement over field genotype
with one of the highest PAR interceptions - 48.72 MWh PAR acre-1 day-1. Fitness trajectories stabilized after approximately 80 generations, indicating convergence of the evolutionary search (Fig. 1F). The similarity among final ideotypes across runs suggests robustness and repeatability of the optimization process.
The following sub-sections characterize the structural and functional features of these optimized architectures/ideotypes, highlighting how architectural modifications improved light distribution throughout the canopy and their broader implications for breeding and crop design.
@@
Daily PAR interception profiles revealed that optimized canopy architectures consistently outperformed the baseline across all hours of the day (Fig. 1D). Improvements were greatest during late afternoon when conventional canopies suffer from self‐shading; the optimized architecture maintained higher PAR penetration throughout the canopy. Simulations across a range of latitudes (SI, Section S4) showed similar gains, indicating that the ideotype is not over‐fit to a single environment. We further evaluated the optimization across a range of planting densities and row spacings; detailed results are provided in SI, Section S5. Radiative‐transfer configuration followed a documented “Helios” framework (reverse ray tracing with multiple scattering), previously shown to agree with field PAR interception measurements, providing additional confidence in the simulated gains [bailey2019helios, saleem2025accessing].
@@
The optimized ideotype featured clear vertical stratification of leaf traits. Leaf angles were measured relative to the stalk, such that smaller angles indicate more erect leaves (0 = parallel to the stalk, 90 = horizontal). Leaves at the top exhibited smaller vertical angles (20–50), effectively minimizing top-layer shading, whereas lower canopy leaves displayed higher vertical angles (60–80), improving light interception in shaded regions. A similar vertical gradient was observed in leaf width, where upper leaves were narrower and lower leaves became progressively wider, enabling efficient capture of the residual sunlight that penetrates deeper into the canopy (Fig. 3). This configuration promotes efficient light redistribution by allowing incoming radiation to penetrate deeper into the canopy and be absorbed more uniformly across layers. Such trait stratification emerged consistently across independently evolved solutions, suggesting a convergent optimization strategy driven by canopy-scale light dynamics.
@@
We also examined how leaves were azimuthally arranged around the stalk, a trait that influences both self-shading and interplant light competition. Maize naturally exhibits an alternate (distichous) phyllotaxy, with leaves emerging in two opposing ranks along the stem. In field-grown maize, this inherent pattern typically manifests as one of three canopy-level azimuthal orientations (see Fig S2): off-row parallel (perpendicular to the planting row), which minimizes mutual shading; on-row parallel, where leaves align with the row direction; or random, where entire individual plants vary in their overall orientation relative to the planting rows [saleem2025accessing, zhou2024genetic].
In contrast to the strict alternate phyllotaxy typical of maize, our EA optimization converged upon a tiled or quasi-radial arrangement of leaf azimuths around the stalk (Fig. 4). This deviation suggests that while the canonical distichous structure may optimize light capture for isolated plants, it generates excessive mutual shading in dense monocultures. The emergent radial tiling represents a ‘community-optimal’ strategy that minimizes self- and neighbor competition.
@@
The optimized ideotype exhibited architectural features consistent with both experimentally validated and commercially successful maize hybrids, underscoring the practical relevance of computational design. Our results on leaf angles closely resemble the smart-canopy architecture characterized by upright upper leaves, moderately inclined middle, and flatter lower leaves—traits that improve light penetration and, consequently, boost grain yield  [ort2015redesigning, mantilla2017differential, tian2024maize]. The resemblance between our computationally derived canopy and the experimentally validated lac1-based F1 hybrids reported by Tian et al. [tian2024maize], which exhibit this smart-canopy configuration, reinforces the physiological credibility and predictive capacity of the optimization framework. A similar pattern occurs in elite commercial hybrids (see, for instance [PioneerLeafAngle2023]), where erect upper leaves enhance light transmission and density tolerance [elli2023maize](Table 1).
@@
To verify that these architectural patterns were mechanistically responsible for improved light capture, we performed trait-restricted optimization experiments in which the EA was restricted to optimize only a subset of parameters while all others were clamped to baseline values. When only leaf angles were free, the algorithm consistently produced ideotypes with erect upper leaves, achieving a modest gain in PAR interception (40.72 $\rightarrow$ 43.00 MWh acre-1 day-1). Optimization of azimuthal angle alone yielded a tiled leaf orientation around the stalk, with a comparable improvement (40.72 $\rightarrow$ 42.99 MWh acre-1 day-1). Allowing angles and widths together led to erect, narrower upper leaves, substantially increasing canopy interception (40.72 $\rightarrow$ 49.47 MWh acre-1 day-1). Finally, when all traits were optimized jointly, the algorithm integrated these features synergistically to reach the highest performance (56.49 MWh acre-1 day-1).
These trait-restricted analyses demonstrate that the key architectural signatures identified in the full optimization (erect angles, radially tiled azimuths, and thinner upper leaves) emerge independently when traits are optimized in isolation, and that their combination yields additive and synergistic benefits to canopy light capture (Table 1).